OpenStudy (anonymous):

One ordered triple (x,y,z) which satisfies the system is (0, 0, 0). What is the only numerical value of "k" for which this system has a solution other than (0,0,0)? x-2y+kz=0 2x-y+3z=0 3x-3y-2z=0

OpenStudy (anonymous):

Let \[A=\left[\begin{matrix}1&-2&k\\ 2&-1&3\\ 3&-3&-2\end{matrix}\right]\] \[A\left[\begin{matrix}x\\y\\z\end{matrix}\right]=\left[\begin{matrix}0\\0\\0\end{matrix}\right]\\ (A^{-1}\cdot A)\left[\begin{matrix}x\\y\\z\end{matrix}\right]=A^{-1}\left[\begin{matrix}0\\0\\0\end{matrix}\right]\\ \left[\begin{matrix}x\\y\\z\end{matrix}\right]=A^{-1}\left[\begin{matrix}0\\0\\0\end{matrix}\right]\] Do you know how to find the inverse of a 3x3 matrix? Also, you might need to know the inverse exists if the determinant of A is non-zero.

OpenStudy (anonymous):

I haven't learn matrix yet

OpenStudy (anonymous):

Is there an easier approach to these problems?

OpenStudy (anonymous):

I think the matrix approach is easier, but that's just my opinion. I'm pretty sure there's something you can do using what you learn about elimination when solving a system of equations. Just a sec...

OpenStudy (anonymous):

\[\begin{cases}x-2y+kz=0\\ 2x-y+3z=0\\ 3x-3y-2z=0\end{cases}\] Multiply the first equation by -2 and add it to the second to eliminate x. Then, multiply the first equation by -3 and add it to the third to eliminate x. \[\begin{cases}-2x+4y-2kz=0\\ 2x-y+3z=0\end{cases}\\ 3y+(3-2k)z=0\ldots(A)\] \[\begin{cases}-3x+6y-3kz=0\\ 3x-3y-2z=0\end{cases}\\ 3y-(2+3k)z=0\ldots(B)\] Rewriting A and B, you get \[3y = -(3-2k)z\\ 3y=(2k-3)z\\\text{ and}\\ 3y=(2+3k)z\] Setting the equations equal to each other, you have \[(2k-3)z=(2+3k)z\\ 2k-3=2+3k\\ k=-5\] So now you just have to solve the system \[\begin{cases}x-2y-5z=0\\ 2x-y+3z=0\\ 3x-3y-2z=0\end{cases}\] Think you can handle that? If I'm not mistaken, I think you should get "many solutions".

OpenStudy (anonymous):

The variables will be cancelled out... x-2y-5z=0 2x-y+3z=0 3x-3y-2z=0 6x-12y-30z=0 6x-3y+9z=0 6x-6y-4z=0 Subtract the first two: -9y-39z=0 subtract the second and third: 3y+13z=0 9y+39z=0 Add those two: (-9y-39z=0) + (9y+39z=0) = 0

OpenStudy (anonymous):

Looks like it's all right. However, I think the last step takes it a bit too far. Right before that, you have \[3y+13z=0, \text{ or}\\ y=-\frac{13}{3}z\] If you plug that into the original first equation, you get \[x=-3z\] This means that if you let z be any (non-zero) number, x and y will also be non-zero because they both depend on z. You just have to be able to write x, y, and z in terms of another variable.

OpenStudy (anonymous):

Oh, I see. Thank you so much!!